OpenStudy (anonymous):

What is the exponent for x when the radical ^3√x^18 is simplified completely?

OpenStudy (debbieg):

Is that cube root? Remember - you can convert radicals to fractional exponents. The power is in the numerator, the root is in the denominator. So \[\Huge \sqrt[3]{x^n}=x^\frac{ n }{3 }\]

OpenStudy (anonymous):

yes its cubed

OpenStudy (debbieg):

OK, see above. Make sense? what do you get for the answer?